OpenStudy (anonymous):

Help me please with (a^4 b^-5) ^-4 I've been trying to get this for an hour and I can't do it

OpenStudy (anonymous):

what do you have to do simplify it?

OpenStudy (anonymous):

I'm not sure it just says that my answer should only contain positive exponents

OpenStudy (anonymous):

oh alright... so you just gotta make it into a fraction... when you multiply the exponents remember the signes will change then negative exponents go into the denominator.. try it and tell me what you got and ill check it\

OpenStudy (anonymous):

I do not know how to make it into a fraction

OpenStudy (jdoe0001):

\(\bf \Large a^{-n} = \cfrac{1}{a^n}\)

OpenStudy (anonymous):

you just have to put the negative exponents in the denominator... i.e x^5v^-5= (x^5)/(v^5)

OpenStudy (anonymous):

What is N?

OpenStudy (anonymous):

it dnsnt rlly matter what it is just that it is negative and he moved it into the denominator

OpenStudy (anonymous):

so (a^4/b^5) ^-4 ???

OpenStudy (anonymous):

all you have to do is move the whole () down think of it as this.. \[\frac{(a^4b^{-5})^{-4}}{1}= \frac{1}{(a^4b^{-5})^4}\] notice atfer we fliped the () down we got a positive 4

OpenStudy (jdoe0001):

\(\bf \Large a^{-cheeseburger} = \cfrac{1}{a^{cheeseburger}}\)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Why didn't the exponets of a abd b change when you put it in the denominator

OpenStudy (anonymous):

*and

OpenStudy (anonymous):

then you gotta distribute your exponent and move some more stuff around\[\frac{ 1 }{(a ^{8}b ^{-20}) }\] then you gotta move b upp and it becomes positive

OpenStudy (anonymous):

could u answer my last question first? I don't get that part

OpenStudy (anonymous):

a^16 not 8 haha

OpenStudy (anonymous):